# MBIAS: Mitigating Bias in Large Language Models While Retaining Context

### v2

@@ 5.2 Performance of MBIAS across Different Demographics
Table 4 shows an analysis of MBIAS’s performance across various demographic groups on the ToxiGen dataset. The key findings are:
@@
Answer Relevancy: The Native American group once again shows the highest Rel. score, at 94.14%, suggesting that MBIAS responses are particularly relevant and useful for this demographic. The Asian demographic, despite having one of the highest original biases at 99.19%, shows sufficient control in post-safety interventions. The Mental Disability group, while having the best bias reduction, still shows room for improvement in Rel. and KR, which are just above 80%. Areas for concern include the Chinese and LGBTQ+ demographics, where higher levels of bias and toxicity persist compared to others. This may be attributed to the data size of these demographics in the test set.
@@
Scores: CN: 4.8, RI: 4.8, CR: 4.8, Len: 4.8
@@
Scores: CN: 4.7, RI: 4.3, CR: 4.4, Len: 4.8
@@
Scores: CN; 4.5, RI: 4.7, CR: 4.5, Len: 4.8
@@
Scores: CN: 4.5, RI: 4.7, CR: 4.8, Len: 4.4
@@
Scores: CN: 4.4, RI: 4.2, CR: 4.3, Len: 4.2
@@
Scores: CN: 4.9, RI: 5.0, CR; 4.9, Len: 4.0
Table 5: Evaluation of MBIAS on benign generation. The table compares original unsafe (biased) statements with their benign counterparts, maintaining the original context. Evaluators scored these transformations on a Likert scale (1-5) across four metrics: Content Neutrality (CN), Respectful Interaction (RI), Content Retention (CR), and Output Length (Len). Higher the score ($\uparrow$), the better. The scores for each output is given below that row.
@@ 5.3 Human Evaluation
Table 5 presents a comparative analysis of biased statements and their benign versions produced by MBIAS. The results show that MBIAS is highly effective in enhancing respectful interaction (RI) and content neutrality (CN), with these categories frequently receiving the highest average scores among the evaluators. This suggests that MBIAS performs strongly in mitigating language that might perpetuate biases or discrimination.
Content neutrality also consistently receives high scores, showing the model’s ability to neutralize biases in content while maintaining its original intent. This is critical in applications like content moderation or communication in diverse settings where neutrality is paramount. The slightly lower scores in output length (Len) might reflect the model’s adaptations in phrasing to achieve neutrality and respectfulness, occasionally necessitating longer or shorter responses than the original. Overall, the data in Table 5 suggests that MBIAS shows great performance in producing outputs that are neutral, respectful, and contextually relevant, enhancing the appropriateness of interactions.
@@ 6 Conclusion
MBIAS is based on the Mistral2-7B-instruct architecture, leveraging instruction-based fine-tuning on a custom dataset designed for safety interventions. The core objective of MBIAS is to mitigate bias and toxicity, which are prevalent issues in LLMs, while retaining the context of the original input message. By embedding debiased or safe alternatives directly within our training dataset, MBIAS effectively recognizing and adjusting biases, ensuring more equitable and balanced content generation. Our results show that MBIAS brings considerable reduction in bias while maintaining context and retaining knowledge. Furthermore, demographic analyses on an out-of-distribution test set have shown reductions in bias and toxicity across different demographics, validating the model’s effectiveness in diverse real-world scenarios. We make the dataset and the model, MBIAS, available to the research community and we encourage further exploration and development in this field.

### v3

@@ 5.2 Performance of MBIAS across Different Demographics
Table 4 shows an analysis of MBIAS performance across various demographic groups on the ToxiGen dataset. The key findings are:
@@
Answer Relevancy: The Native American group once again shows the highest Rel. score, at 94.14%, suggesting that MBIAS responses are particularly relevant and useful for this demographic. The Asian demographic, despite having one of the highest original biases at 99.19%, shows sufficient control in post-safety interventions. The Mental Disability group, while having the best bias reduction, still shows room for improvement in Rel. and KR, which are just above 80%. There are still many areas for concern, including the Women, LGBTQ+, Chinese, Native American, Middle Eastern, Mexican, and Latino demographics, where higher levels of bias (above 15%) and toxicity persist compared to others.
Finding: MBIAS effectively debiases text across various demographics. Certain demographics, such as Mental Disability, benefit the most from MBIAS, with the highest bias reduction, while Native American experiences the most significant toxicity reduction. The minor discrepancies in scores may be attributed to the distribution of the data samples across different demographics.
@@
Scores: Content Neutrality: 4.8, Respectful Interaction: 4.8, Content Retention: 4.8, Output Length: 4.8
@@
Scores: Content Neutrality: 4.7, Respectful Interaction: 4.3, Content Retention: 4.4, Output Length: 4.8
@@
Scores: Content Neutrality: 4.5, Respectful Interaction: 4.7, Content Retention: 4.5, Output Length: 4.8
@@
Scores: Content Neutrality: 4.5, Respectful Interaction: 4.7, Content Retention: 4.8, Output Length: 4.4
@@
Scores: Content Neutrality: 4.4, Respectful Interaction: 4.2, Content Retention: 4.3, Output Length: 4.2
@@
Scores: Content Neutrality: 4.9, Respectful Interaction: 5.0, Content Retention: 4.9, Output Length: 4.0
Table 5: Evaluation of MBIAS on benign generation. The table compares original unsafe (biased) statements with their benign counterparts, maintaining the original context. Evaluators scored these transformations on a Likert scale (1-5) across four metrics: Content Neutrality, Respectful Interaction, Content Retention, and Output Length. Higher the score ($\uparrow$), the better. The scores for each output are given below the corresponding row.
@@ 5.3 Human Evaluation
Table 5 presents a comparative analysis of biased statements and their benign versions produced by MBIAS. The results show that MBIAS is highly effective in enhancing respectful interaction and content neutrality , with these categories frequently receiving the highest average scores among the evaluators. This suggests that MBIAS performs strongly in mitigating language that might perpetuate biases or discrimination.
Content neutrality also consistently receives high scores, showing the model ability to neutralize biases in content while maintaining its original intent. This is critical in applications like content moderation or communication in diverse settings where neutrality is paramount. The slightly lower scores in output length might reflect the model adaptations in language generation to achieve neutrality and respectfulness, occasionally necessitating longer or shorter responses than the original.
Finding Overall, the results in Table 5 suggests that our instruction fine-tuning method MBIAS shows great performance in producing outputs that are neutral, respectful, and contextually relevant, enhancing the appropriateness of interactions.
5.4 Error Analysis
The error analysis of our method, which has undergone safety fine-tuning, reveals several key insights into its performance and limitations. Despite rigorous fine-tuning, the model can still occasionally generate outputs that can be deemed unsafe or inappropriate. One common error type is the occasional generation of biased or offensive language, especially in contexts involving sensitive topics such as race, gender, or religion. This indicates that while the fine-tuning process has significantly reduced the frequency of such outputs, it has not entirely eliminated them.
Another potential error involves the model tendency to provide misinformation or factually incorrect statements. This highlights the challenge of ensuring accuracy in language models, as fine-tuning for safety does not inherently guarantee factual correctness. Additionally, the model sometimes produces contextually inappropriate responses, which can be attributed to the complexities of human language and the diverse ways in which prompts can be interpreted. These errors underscore the need for continuous monitoring and iterative improvement in the fine-tuning process to enhance the model safety and reliability further.
While we utilize LLM as evaluator, it is important to acknowledge that they may inherit biases from their training data, which can skew their judgments. These biases can manifest in various forms, including systematic bias, where certain demographics or viewpoints are unfairly represented or evaluated. To mitigate this, we complement LLM evaluation with human assessment, albeit on a smaller subset.
Future work should focus on developing more sophisticated fine-tuning techniques and incorporating real-time feedbacks. Enhancements LLM as judge should prioritize statistical methods to ensure fair and unbiased evaluations.
@@ 6 Conclusion
MBIAS is built on the top of Mistral2-7B-instruct architecture, leveraging instruction-based fine-tuning on a custom dataset designed for safety interventions. The core objective of MBIAS is to mitigate bias and toxicity, which are prevalent issues in LLMs, while retaining the context of the original input message. By embedding debiased or safe alternatives directly within our training dataset, MBIAS effectively recognizing and adjusting biases, ensuring more equitable and balanced content generation. Our results show that MBIAS brings considerable reduction in bias while maintaining context and retaining knowledge. Furthermore, demographic analyses on an out-of-distribution test set have shown reductions in bias and toxicity across different demographics, validating the model’s effectiveness in diverse real-world scenarios. We make the dataset and the model, MBIAS, available to the research community for reproducibility and further research.
@@
Acknowledgements
We extend our gratitude to the Province of Ontario, the Government of Canada through CIFAR, and the corporate sponsors of the Vector Institute for their generous support and provision of resources essential for this research. Further details on our sponsors can be found at www.vectorinstitute.ai/#partners. We also acknowledge our expert review team and everyone involved in the data review process.